Midland Company buys tiles and prints different designs on them for souvenir and gift stores. It buys the tiles from a small company in Europe, so at all times it keeps on hand a stock equal to the tiles needed for three months’ sales. The tiles cost $1.00 each and must be paid for in cash. The company has 27,000 tiles in stock. Sales estimates, based on contracts received, are as follows for the next six months:

January 13,100

February 17,000

March 13,100

April 14,000

May 9,500

June 8,000

Required: a. & b. Estimate purchases (in units) and cash required to make purchases in January, February, and March.
Answer #1

Answer-a&b-

MIDLAND COMPANY
MERCHANDISE PURCHASE BUDGET
FOR THE PERIOD ENDED MARCH 31
PARTICULARS JANUARY FEBRUARY MARCH
Estimated unit sales 13100 17000 13100
Add- Estimated sales for next three months 17000+13100+14000= 44100 13100+14000+9500= 36600 14000+9500+8000= 31500
Total inventory needs 57200 53600 44600
Less- Beginning inventory 27000 44100 36600
Units to be purchased (a) 30200 9500 8000
Unit cost (b) $1.00 $1.00 $1.00
Estimated purchase cost (c=a*b) $30200 $9500 $8000